OpenStudy (anonymous):

x^-2 - y^-2 _____________ x^-1 - y^-1

OpenStudy (anonymous):

there is a slick way to do this, but it is more meaningful if you write what this really is without negative exponents

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{\frac{1}{x^2}-\frac{1}{y^2}}{\frac{1}{x}-\frac{1}{y}}\]

OpenStudy (anonymous):

ok then i find the lcd of the top and bottom

OpenStudy (anonymous):

now multiply top and bottom by \[x^2y^2\] to clear the fractions

OpenStudy (anonymous):

lcd

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

right. the slick way is to write \[\frac{1}{x^2}-\frac{1}{y^2}=(\frac{1}{x}+\frac{1}{y})(\frac{1}{x}+\frac{1}{y})\] and then cancel

OpenStudy (anonymous):

difference of squares

OpenStudy (anonymous):

typo there should be \[\frac{1}{x^2}-\frac{1}{y^2}=(\frac{1}{x}-\frac{1}{y})(\frac{1}{x}+\frac{1}{y})\]

OpenStudy (anonymous):

cancels the bottom

OpenStudy (anonymous):

yes if you do it that way and cancel, you get \[\frac{1}{x}+\frac{1}{y}\] in one step

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

yes unless you care to add them and get \[\frac{x+y}{xy}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks bro